OpenStudy (mik812):

• 21) Find the unit-less area of the ellipse with foci (-2, 0), (2, 0) and y-intercepts -2 and 2.

OpenStudy (irishboy123):

you know that \(f^2 = a^2 - b^2\) and \(A = \pi ab\) it's all here https://en.wikipedia.org/wiki/Ellipse

likeabossssssss (likeabossssssss):

welcome to os